# Towards a Faithful Quantumness Certification Functional for One-Dimensional Continuous-Variable Systems

### v1

@@ Conclusions
We have derived the most sensitive family of quantumness certification functionals to date. Their transparent form hopefully opens up the door for further improvements eventually resulting in a general and useful faithfully discriminating approach to quantumness certification.
@@ Appendix A Certification functionals ${\tilde{\xi}}(S,k)$
In Ref. [1] slightly generalized two-parameter families of nonclassicality certification functions were introduced, namely, expressions of the form
@@
with $S_{k}=1-\frac{1-S}{k}$, and where the parameter $S\leq 1$ selects the reference distribution ${\tilde{\cal P}}(x,p;S)$, see Eq. (1). Expression (6) can be useful for some field correlation measurements [1].
@@
As a function of $k$, negative minimum values ${\tilde{\xi}}_{-}(S_{k})$ with the greatest magnitude, occur at the balanced mid-point, $k=\frac{1}{2}$, providing the greatest contrast when applying $\tilde{\xi}$. This is why here, without loss of generality, we confine our discussion of $\xi$ to the form of Eq. (2) only.

### v2

@@ Conclusions
We have generalized the quantumness certification functional of Ref. [3]. We have proved that the quantumness functionals (4) always identify classical states correctly, but for weakly nonclassical states they can fail, incorrectly mislabelling these as classical. We proved that these quantumness certification functionals always identify mixtures of states as a concave combination of their constituents. Our generalized certification functionals’ transparent form hopefully opens up the door for further improvements eventually resulting in a general and useful faithfully discriminating approach to quantumness certification. Its transparent form has already allowed us to generalize to multimode settings, see Eq. (5).
@@ Appendix A Certification functionals ${\tilde{\xi}}(S,k)$
In Ref. [3] the two-parameter families of nonclassicality certification functions
@@
with $S_{k}=1-\frac{1-S}{k}$, were introduced. Here the parameter $S\leq 1$ selects the reference distribution ${\tilde{\cal P}}(x,p;S)$, see Eq. (1). Expression (9) can be useful for some field correlation measurements [3].
@@
As a function of $k$, negative minimum values ${\tilde{\xi}}_{-}(S_{k})$ with the greatest magnitude, occur at the balanced mid-point, $k=\frac{1}{2}$, providing the greatest contrast when applying $\tilde{\xi}$ [2]. This is why here, without loss of generality, we confine our discussion of $\xi$ to the form of Eq. (2) only.